word problem delimma tickets for a play cost $12 for an adult and $8 for an child If 130 tickets were sold and the total receipts were 1160 how many of each ticket weree sold. Ok if x= the adult tickets and c= the child tickets do I go x+ c = 130 and then go 12x + 8c==1160 then divide both sidec by 8 to get the child ticket
Answer by rossiv53(27)

Lets begin by assigning a variable to one of the ticket values.
x = number of adult tickets sold
Since there were a total of 130 tickets sold, the amount of children's tickets sold can be expressed as (130-x)
Now you can set up your equation
12x+8(130-x) = 1160
12x+1040-8x = 1160
4x+1040 = 1160
4x+1040-1040 = 1160-1040
4x = 120
x = 30
So, 30 adult tickets were sold and 100 children's tickets were sold
Good luck!